# Topology of shallow-water waves on the rotating sphere

### v1

@@ 3 Modal flow and meridional zeros of eigenfunctions
To this day, Equations (2.1) do not have an exact solution for arbitrary values of $\epsilon$, although good approximations have been found in certain limits, especially in short wavelengths ranges [32, 36], or for special values of $\omega$ and $\epsilon$ [34]. In this study, we are interested with the transition between the Matsuno and Margules limits, i.e. the evolution of eigenmodes and eigenfrequencies for finite values of $\epsilon$. For this reason, the analysis of this section is mostly based on numerical integration of Equations (2.1), with the spectral solver Dedalus [37]. The point of this section is the following: the Matsuno spectrum, i.e. the frequencies $\omega$ of zonally-propagating equatorial waves on the unbounded $\beta$-plane, consists in a discrete set of continuous functions of the zonal wave number $k$, or spectral branches. These branches can be indexed by the number of zeros of the meridional velocity $\tilde{v}(\theta)$. For shallow-water waves on the sphere, however, the eigenmodes and eigenfrequencies depart from Matsuno’s solutions as $\epsilon$ increases, eigenfunctions spread away from the equator and new zeros of $\tilde{v}(\theta)$ appear at non-zero latitudes, modifying the natural order of mode branches established by Matsuno. This affects the conclusion regarding the very concept of spectral flow in this model, which must be replaced by the more accurate notion of modal flow.
3.1 Reminder: the Matsuno spectrum and eigenfunctions
@@
The shallow-water equations for equatorial waves on the unbounded $\beta$-plane are the same as (2.1), only with Cartesian coordinates instead of the spherical ones ($x$ is the zonal coordinate pointing eastward and $y$ is the meridional coordinate pointing northward, with $y=0$ defining the equator), and $f=\beta y$. Once again, we focus on plane waves propagating in the zonal direction, i.e. solutions in the form $X(y){\rm e}^{{\rm i}(kx-\omega t)}$, where the wave number $k$ can take continuous values. As explained in the introduction, this model captures the physics of equatorial waves in the limit of small $\epsilon$. The solutions found in [13] obey the dispersion relation
@@
and the corresponding eigenfunctions for the meridional velocity are given by
@@
Figure 5: The Matsuno spectrum and eigenfunctions for equatorial shallow-water waves on the unbounded $\beta$-plane [13]. Left: plot of the dispersion relation (8), showing the branches $-1\leq p\leq 3$ for positive frequencies. The Yanai and Kelvin branches transit across the frequency gap between the Rossby and inertia-gravity wavebands, as $k$ increases. Right: Amplitude of the meridional velocity in the $y$ direction, given by Expression (9). The branch index $p$ is equal to the number of zeros of $\tilde{v}(y)$.
@@ 3.2 Zeros of the meridional velocity
The $+2$ spectral flow studied in [1] corresponds to the Yanai and Kelvin modes, which transit from the Rossby waveband to the inertia-gravity waveband as $k$ increases. Beyond the fact that these modes form continuous curves in $(k,\omega)$, what identifies each of them is the number of zeros of the meridional velocity in the $y$ direction, as previously shown. Generally speaking, it is straightforward to define the spectral flow of an eigenvalue problem if there is a continuous parameter such as the zonal wave number $k$, because the spectrum thus consists in continuous branches. However, if the spectral parameter takes discrete values, one may rather identify modes of the same branch with their number of zeros for a given field. For shallow-water waves on the rotating sphere, since the azimuthal wave number $m$ is a discrete parameter, counting the number of zeros of $\tilde{v}(\theta)$ on the open interval $(-\pi/2,+\pi/2)$ thus seems to be a natural way of ordering the eigenmodes of the problem (2.1) (see Figure 6). In fact, many works on the spectrum of shallow-water waves associate or label modes according to the number of zeros of their meridional velocity [31, 32, 59, 34, 43]. In the Matsuno spectrum, at negative $k$, there are modes with no zero of $\tilde{v}$ in the Rossby waveband but not in the inertia-gravity waveband, whereas there are modes with no zero in the inertia-gravity waveband at positive $k$. Connecting them together forms the spectral flow of Yanai modes, which transits between the Rossby and inertia-gravity wavebands as $k$ increases. To clarify, in the rest of the paper, we will employ the term modal branch to refer to modes with the same number of zeros of $\tilde{v}(\theta)$ on the open interval $(-\pi/2,\pi/2)$, as $m$ varies, and say that there is a modal flow when such a modal branch transits between the Rossby and inertia-gravity wavebands. In contrast, the terms spectral branch and spectral flow will be used to refer to the continuous curves formed by the eigenfrequencies, which is a well-defined concept only if the spectral parameter is continuous, strictly speaking. Spectral and modal branches of the Matsuno spectrum are the same (see Figure 5). In the shallow-water spectrum on the sphere, the spectral parameter $m$ is not continuous, yet one can argue that there is a spectral flow of eigenfrequencies in the Matsuno limit (see Figures 1 and 4), but not in the Margules limit.
@@
Figure 6: Numerically-calculated eigenvalues of Equations (2.1) (red dots). The modes with same number of zeros of $\tilde{v}(\theta)$ are connected by dashed blue curves, which have no physical meaning. The number of zeros is indicated on the curves. For small $\epsilon$, the modal branches display an abrupt jump right before $m=0$. As $\epsilon$ increases, the spectrum progressively loses its east-west asymmetry and the spectral flow collapses.
@@
As shown in the previous section, for small $\epsilon$, the ratio between the equatorial radius of deformation and the sphere’s radius is of order $\sqrt{\epsilon}$, which implies that the eigenfunctions spread across the whole sphere even for moderate values of $\epsilon$, and thus the eigenmodes experience the discrepancy with the unbounded $\beta$-plane. The spreading of all the eigenfunctions reveals additional zeros of $\tilde{v}$ at non-zero latitude, which are not visible when $\epsilon$ is too small and the eigenmodes are strongly trapped at the equator. In fact, all modes with positive phase speed (i.e. positive $m$, considering only the positive frequencies) have two additional zeros at opposite latitudes (see Figures 7 and 8), compared to the same modes of the unbounded $\beta$-plane. However, the zeros of Rossby modes, which propagate westward, are unchanged. In other words, the index $p$ of the Matsuno modes propagating eastward is increased by $+2$ in spherical geometry. In particular, Yanai modes with negative $m$ have $p=0$ meridional zeros and $p=2$ zeros for positive $m$. Similarly, in contrast with the unbounded $\beta$-plane, Kelvin modes have a non-zero meridional velocity which cancels once at the equator, i.e. $p=1=-1+2$. This $+2$ jump in number of zeros, which was already discussed by Müller and O’Brien in [34], is illustrated in Figure 6.
@@
Figure 7: Meridional velocity $\mathrm{Re}\left(\tilde{v}(\theta){\rm e}^{{\rm i}m\phi}\right)$ of eigenmodes for $\epsilon=0.2$ and $m=-2,+1$, computed with Dedalus (red: positive values; blue: negative values). A is a Rossby mode, B and F are Yanai modes, E is a Kelvin mode, C, D, G and H are inertia-gravity modes. On the left, modes are connected by dashed blue curves according to number of zeros of $\tilde{v}(\theta)$ on $(-\pi/2,\pi/2)$. Modes F,G and H have zeros at high latitude, which are not visible on the plot since the eigenfunctions vanish away from the equator (see Appendix 6.1). Dashed black lines indicate the zeros of $\tilde{v}(\theta)$, and the meridional lines of zeros correspond to the term ${\rm e}^{{\rm i}m\phi}$ with $m\neq 0$ for travelling waves.
@@
Figure 8: Same modes as Figure 7, for $\epsilon=2$. The eigenfunctions are more spread in latitude and extend up to the poles, clearly revealing zeros at high latitudes. The meridional velocity of a Yanai mode has no zero for negative $m$ (B), but changes sign at two opposite latitudes for positive $m$ (F), contrary to the Yanai mode computed by Matsuno on the equatorial $\beta$-plane. Generally speaking, the meridional velocity of modes with eastward phase speed have two more zeros, compared to Matsuno’s eigenfunctions. Similarly, Kelvin modes (E) have zero meridional velocity on the unbounded $\beta$-plane, which is not true on the sphere.
@@
To conclude this part, we have shown that the net modal flow of shallow-water waves, equal to $+2$ on the unbounded $\beta$-plane, is null on the sphere. This is evident for large $\epsilon$ since the spectrum displays a strong east-west symmetry, i.e. the eigenfrequencies $\omega$ and eigenmodes are nearly the same for $m$ and $-m$. As $\epsilon$ becomes smaller, the spectrum loses this east-west symmetry, displays a spectral flow of eigenvalues but the net modal flow remains zero. At small $\epsilon$, the spectrum resembles the Matsuno one as there are Kelvin and Yanai modes transiting across the gap for increasing wave number $m$. However, the Yanai modes do not form a modal branch because the eastward ones have two more zeros than the westward ones. The modal flow correctly tells how many modes are gained by a waveband as $m$ increases, and this number is sometimes referred to as analytical index [16]. In the present case, its nullity can be surprising regarding the bulk-interface correspondence established in [1, 60, 16], which states that the number of transiting modes of the Matsuno spectrum is equal to a robust topological charge $+2$ at the equator. In the following we will show that the nullity of the net modal flow can actually be predicted with the index theorem, as the Chern numbers of the shallow-water spectrum on the sphere add up to zero, in contrast with the unbounded $\beta$-plane.
@@ 4 Topology of the shallow-water model on the sphere
Identifying modes with the number of zeros of their $\tilde{v}$ eigenfunction, we showed that the net modal flow of shallow-water waves on the sphere is null, while it is $+2$ for the unbounded $\beta$-plane. Many studies [61, 1, 18, 22, 19, 62, 20, 21, 16] showed that, when the frequencies $\omega$ can be expressed as the eigenvalues of a Hermitian differential wave operator $\mathcal{H}$, the number of modes gained by a waveband as the spectral parameter increases is equal to a topological invariant characterising the Weyl symbol $H$ of the operator, in virtue of the index theorem [15]. As demonstrated for instance in [1, 15, 16], the presence of $+2$ transiting equatorial modes on the unbounded $\beta$-plane is ensured by a $+2$ Chern number characterising the degeneracy point of the Weyl symbol $H$ of the wave operator $\mathcal{H}$ at $f=0$. More precisely, the Yanai and Kelvin branches constitute a footprint, in the spectrum of $\mathcal{H}$, of a degeneracy point of its symbol bearing a Chern number $+2$. In light of the discussion of section 3, the aim of this section is to show that the previous conclusions can be inferred with the index theorem, as accounting for the spherical metric reveals new degeneracy points of the symbol of the wave operator, in addition to the unique degeneracy point of the problem on the unbounded $\beta$-plane. We will demonstrate that these additional degeneracy points bear topological charges which compensate the $+2$ Chern number of the equatorial degeneracy point, thus making the eigenbands of the symbol have zero total Chern numbers, which justifies the absence of transiting modes and the breaking of Yanai and Kelvin mode branches on the sphere for large $\epsilon$.
@@ 4.1 Hermitian form, analogy with topographic shallow-water waves
In order to apply the index theorem, we first need to express Equations (2.1) in the form $\omega X=\mathcal{H}X$, where $\mathcal{H}$ is a Hermitian differential wave operator and the complex eigenvector $X$ contains the three dependent variables $\tilde{u}(\theta),\tilde{v}(\theta)$ and $\tilde{h}(\theta)$. To do this, we introduce the rescaled fields $u=\sqrt{\cos(\theta)}\,\tilde{u}$, $v=\sqrt{\cos(\theta)}\,\tilde{v}$ and $\eta=\sqrt{\cos(\theta)}\,\tilde{h}$. Equations (2.1) can thus be recast in the form of a matrix eigenvalue equation:
$\omega\begin{pmatrix}u\\[6.0pt] v\\[6.0pt] \eta\end{pmatrix}=\begin{pmatrix}0&&{\rm i}f&&\frac{c}{R}\frac{m}{\cos\theta}\\[6.0pt] -{\rm i}f&&0&&-{\rm i}\frac{c}{R}\frac{{\rm d}}{{\rm d}\theta}-{\rm i}\beta_{\rm g}\\[6.0pt] \frac{c}{R}\frac{m}{\cos\theta}&&-{\rm i}\frac{c}{R}\frac{{\rm d}}{{\rm d}\theta}+{\rm i}\beta_{\rm g}&&0\end{pmatrix}\begin{pmatrix}u\\[6.0pt] v\\[6.0pt] \eta\end{pmatrix}=\mathcal{H}\begin{pmatrix}u\\[6.0pt] v\\[6.0pt] \eta\end{pmatrix}\ ,$ (10)
@@
where we have also defined $M=c\>m/(R\cos\theta)$, and $K_{\theta}$ the symbol of the Hermitian differential operator $-{\rm i}{\rm d}/{\rm d}\theta$ times $c/R$, which is therefore a real quantity (see [63], Theorem 13.8). A recent paper [64] wrongly states that the shallow-water wave operator on the sphere is non-Hermitian and admits complex eigenvalues, which is not true since this system is linearly stable in the absence of a background flow. The mistake originates from using the Fourier transform (which is wrong since the coefficients of $\mathcal{H}$ vary with $\theta$) on the wave operator, instead of the more relevant Wigner transform which gives its Weyl symbol. It is known that the Wigner transform must include additional terms on curved manifolds [65], which yields a Hermitian symbol in this case. Moreover, the equivalence between the respective Hermiticity of an operator and its symbol is a known property [63], and the symbol (12) is Hermitian as we used an appropriate rescaling of the variables, implying the Hermiticity of the operator $\mathcal{H}$ for the canonical scalar product (see Appendix 6.2).
@@
$H$ is thus a continuous matrix function of the conjugate variables $\theta,K_{\theta}$, and $m$. To be clear, $m$ must be an integer so that the eigenmodes of $\mathcal{H}$ are regular on the sphere, however we can formally consider the symbol $H$ as a continuous function of $m$ or $M$. In a nutshell, the symbol $H$ represents quasi-local dispersion and polarisation relations of shallow-water waves (see e.g. [66, 67, 68, 69]). In the context of this paper, we introduce it in order to apply the index theorem, following [15], i.e. we investigate the topological properties of the degeneracy points of $H$. This theorem relates the Chern numbers of these degeneracy points to the number and direction of transiting modes of the wave operator $\mathcal{H}$, which can be interpreted as the spectral footprint of a degeneracy point of non-zero Chern number in the quasi-local dispersion relations. The symbol (12) is a generalisation of the one studied in [1] for the unbounded $\beta$-plane: indeed, in the Matsuno limit, $\beta_{g}/f\approx\epsilon/4\ll 1$ in the Yoshida waveguide, thus the metric term $\beta_{g}$ can be dropped, as far as equatorially-trapped waves are concerned. Equations (10) are formally equivalent to the shallow-water model in planar geometry with varying topography, whose transiting modes were investigated in [19]. The metric term $\beta_{\rm g}$ plays the exact same role as the topographic parameter $\beta_{t}$, although [19] only considered an $f$-plane. In other words, the metric $\beta$-term is equivalent to adding topography in flat metric. Actually one could consider both the spherical metric with a topography $h_{0}(\theta)$ varying with latitude (thus the velocity $c=\sqrt{gh_{0}}$ is also a function of latitude). It can be shown that the symbol of this problem is the same as (12), with a total $\beta$-term that is the sum of $\beta_{\rm g}$ and $\beta_{t}$, thus adding up to
$\beta_{\text{total}}=\frac{c}{2R}\left(\tan\theta-\frac{{\rm d}\ln c}{{\rm d}\theta}\right)\ .$ (13)
We will not be considering a varying topography, as we wish to focus on the curved metric. However, we bring to the analysis of [19] the case of varying $f$ and the additional geometric/topographic term with latitude, which was already addressed in part 3.5.1 of [38], in the context of the equatorial channel.

### v2

@@ 3 Modal flow and zeros of the meridional velocity
To this day, Equations (4) do not have an exact solution for arbitrary values of $\epsilon$, although good approximations have been found in certain limits, especially in short wavelength ranges [17, 21], or for special values of $\omega$ and $\epsilon$ [19]. In this study, we are interested in the transition between the Matsuno and Margules limits, i.e. the evolution of modes and frequencies for finite values of $\epsilon$. For this reason, the analysis of this section is mostly based on numerical integration of Equations (4), with the spectral solver Dedalus [22]. The point of this section is the following: the Matsuno spectrum, i.e. the frequencies $\omega$ of zonally-propagating equatorial waves on the unbounded $\beta$-plane, consists of a discrete set of continuous functions of the zonal wave number $k$, or spectral branches. These branches can be indexed by the number of zeros of the meridional velocity $\tilde{v}(\theta)$. For shallow-water waves on the sphere, however, the modes and frequencies depart from Matsuno’s solutions as $\epsilon$ increases, wave functions spread away from the equator and new zeros of $\tilde{v}(\theta)$ appear at non-zero latitudes, modifying the natural order of mode branches established by Matsuno. This affects the conclusion regarding the very concept of spectral flow in this model, which must be replaced by a more accurate notion that we name the modal flow.
3.1 Reminder: the Matsuno spectrum and wave functions
@@
The shallow-water equations for equatorial waves on the unbounded $\beta$-plane are the same as (4), only with Cartesian coordinates instead of the spherical ones ($x$ is the zonal coordinate pointing eastward and $y$ is the meridional coordinate pointing northward, with $y=0$ defining the equator), and $f=\beta y$. Once again, we focus on plane waves propagating in the zonal direction, i.e. solutions in the form $X(y){\rm e}^{{\rm i}(kx-\omega t)}$, where the wave number $k$ can take continuous values. As explained in the introduction, this model captures the physics of equatorial waves in the limit of small $\epsilon$. The solutions found by [13] obey the dispersion relation
@@
and the corresponding wave functions for the meridional velocity are given by
@@
Figure 5: The Matsuno spectrum and wave functions for equatorial shallow-water waves on the unbounded $\beta$-plane [13]. Left: plot of the dispersion relation (9), showing the branches $-1\leq p\leq 3$ for positive frequencies. The Yanai and Kelvin branches transit across the frequency gap between the Rossby and inertia-gravity wavebands, as $k$ increases. Right: Amplitude of the meridional velocity in the $y$ direction, given by Expression (10). The branch index $p$ is equal to the number of zeros of $\tilde{v}(y)$.
@@ 3.2 Zeros of the meridional velocity
The $+2$ spectral flow studied by [1] corresponds to the Yanai and Kelvin modes, which transit from the Rossby waveband to the inertia-gravity waveband as $k$ increases. Beyond the fact that these modes form continuous curves in $(k,\omega)$, what identifies each of them is the number of zeros of the meridional velocity in the $y$ direction, as previously shown. Generally speaking, it is straightforward to define the spectral flow of an eigenvalue problem if there is a continuous parameter such as the zonal wave number $k$, because the spectrum thus consists of continuous branches. However, if the spectral parameter takes discrete values, one must find an alternative way to count the number of modes gained or lost by a waveband as this spectral parameter is swept. For shallow-water waves on the rotating sphere, the azimuthal wave number $m$ is a discrete parameter, nevertheless we will extrapolate the relation existing between the spectral flow of inertia-gravity waves and the number of zeros of $\tilde{v}$ for the unbounded $\beta$-plane model. In other words, we will adopt the number of zeros of $\tilde{v}(\theta)$ on the open interval $(-\pi/2,+\pi/2)$ as an ordering parameter for the modes of the problem (4) (see Figure 6). In fact, many works on the spectrum of shallow-water waves associate or label modes according to the number of zeros of their meridional velocity [16, 17, 58, 19, 42]. In the Matsuno spectrum, at negative $k$, there are modes with no zero of $\tilde{v}$ in the Rossby waveband but not in the inertia-gravity waveband, whereas there are modes with no zero in the inertia-gravity waveband at positive $k$. Connecting them together forms the spectral flow of Yanai modes, which transits between the Rossby and inertia-gravity wavebands as $k$ increases. To clarify, in the rest of the paper, we will employ the term modal branch to refer to modes with the same number of zeros of $\tilde{v}(\theta)$ on the open interval $(-\pi/2,\pi/2)$, as $m$ varies, and say that there is a modal flow when such a modal branch transits between the Rossby and inertia-gravity wavebands. In contrast, the terms spectral branch and spectral flow will be used to refer to the continuous curves formed by the frequencies, which is a well-defined concept only if the spectral parameter is continuous, strictly speaking. Spectral and modal branches of the Matsuno spectrum are the same (see Figure 5). In the shallow-water spectrum on the sphere, the spectral parameter $m$ is not continuous, yet one can argue that there is a spectral flow of frequencies in the Matsuno limit (see Figures 1 and 4), but not in the Margules limit.
@@
Figure 6: Numerically-calculated frequencies of Equations (4) (red dots). The modes with same number of zeros of $\tilde{v}(\theta)$ are connected by dashed blue curves, which have no physical meaning. The number of zeros is indicated on the curves. For small $\epsilon$, the modal branches display an abrupt jump right before $m=0$. As $\epsilon$ increases, the spectrum progressively loses its east-west asymmetry and the spectral flow collapses.
@@
As shown in the previous section, for small $\epsilon$, the ratio between the equatorial radius of deformation and the sphere’s radius is of order $\sqrt{\epsilon}$, which implies that the wave functions spread across the whole sphere even for moderate values of $\epsilon$, and thus the modes experience the discrepancy with the unbounded $\beta$-plane. The spreading of all the wave functions reveals additional zeros of $\tilde{v}$ at non-zero latitude, which are virtually invisible when $\epsilon$ is too small and the modes are strongly trapped at the equator. In fact, all modes with positive phase speed (i.e. positive $m$, considering only the positive frequencies) have two additional zeros at opposite latitudes (see Figures 7 and 8), compared to the same modes of the unbounded $\beta$-plane. However, the zeros of Rossby modes, which propagate westward, are unchanged. In other words, the index $p$ of the Matsuno modes propagating eastward is increased by $+2$ in spherical geometry. In particular, Yanai modes with negative $m$ have $p=0$ meridional zeros and $p=2$ zeros for positive $m$. Similarly, in contrast with the unbounded $\beta$-plane, Kelvin modes have a non-zero meridional velocity which cancels once at the equator, i.e. $p=1=-1+2$. This $+2$ jump in number of zeros, which was already discussed by [19], is illustrated in Figure 6.
@@
Figure 7: Meridional velocity $\mathrm{Re}\left(\tilde{v}(\theta){\rm e}^{{\rm i}m\phi}\right)$ of modes for $\epsilon=0.2$ and $m=-2,+1$, computed with Dedalus (red: positive values; blue: negative values). A is a Rossby mode, B and F are Yanai modes, E is a Kelvin mode, C, D, G and H are inertia-gravity modes. On the left, modes are connected by dashed blue curves according to number of zeros of $\tilde{v}(\theta)$ on $(-\pi/2,\pi/2)$. Modes F,G and H have zeros at high latitude, which are not visible on the plot since the wave functions vanish away from the equator (see Appendix 6.1). Dashed black lines indicate the zeros of $\tilde{v}(\theta)$, and the meridional lines of zeros correspond to the term ${\rm e}^{{\rm i}m\phi}$ with $m\neq 0$ for travelling waves.
@@
Figure 8: Same modes as Figure 7, for $\epsilon=2$. The wave functions are more spread in latitude and extend up to the poles, clearly revealing zeros at high latitudes. The meridional velocity of a Yanai mode has no zero for negative $m$ (B), but changes sign at two opposite latitudes for positive $m$ (F), contrary to the Yanai mode computed by Matsuno on the equatorial $\beta$-plane. Generally speaking, the meridional velocity of modes with eastward phase speed has two more zeros, compared to Matsuno’s wave functions. Similarly, Kelvin modes (E) have zero meridional velocity on the unbounded $\beta$-plane, which is not true on the sphere. Note that the north-south symmetry of the wave functions is preserved on the sphere, which implies the parity of these additional zeros. This point is discussed in Appendix 6.2.
@@
To conclude this part, we have shown that the net modal flow of shallow-water waves, equal to $+2$ on the unbounded $\beta$-plane, is null on the sphere. This is evident for large $\epsilon$ since the spectrum displays a strong east-west symmetry, i.e. the frequencies $\omega$ and wave functions are nearly the same for $m$ and $-m$. As $\epsilon$ becomes smaller, the spectrum loses this east-west symmetry, displays an apparent spectral flow of frequencies but the net modal flow remains zero. At small $\epsilon$, the spectrum resembles the Matsuno one as there are Kelvin and Yanai modes transiting across the gap for increasing wave number $m$. However, the Yanai modes do not form a modal branch because the eastward ones have two more zeros than the westward ones. In the present case, the nullity of this modal flow can be surprising regarding the bulk-interface correspondence established by [1, 59, 24], which states that the number of transiting modes of the Matsuno spectrum is equal to a robust topological charge $+2$ at the equator. In the following we will show that the nullity of the net modal flow is actually in agreement with the index theorem, as the Chern numbers of the shallow-water spectrum on the sphere add up to zero, in contrast with the unbounded $\beta$-plane.
@@ 4 Topology of the shallow-water model on the sphere
Identifying modes with the number of zeros of their $\tilde{v}$ wave function, we showed that the net modal flow of shallow-water waves on the sphere is null, while it is $+2$ for the unbounded $\beta$-plane. Many studies [60, 1, 26, 30, 27, 61, 28, 29, 24] showed that, when the frequencies $\omega$ can be expressed as the eigenvalues of a Hermitian differential wave operator $\mathcal{H}$, the number of modes gained by a waveband as the spectral parameter increases is equal to a topological invariant characterising the Weyl symbol $H$ of the operator $\mathcal{H}$, in virtue of the index theorem [23]. As demonstrated for instance by [1, 23, 24, 62], the presence of $+2$ transiting equatorial modes on the unbounded $\beta$-plane is ensured by a $+2$ Chern number characterising the degeneracy point of the symbol at $f=0$. More precisely, the Yanai and Kelvin branches constitute a footprint, in the spectrum of $\mathcal{H}$, of a degeneracy point of its symbol bearing a Chern number $+2$. In light of the discussion of section 3, the aim of this section is to show that the previous conclusions can be inferred with the index theorem, as accounting for the spherical metric reveals new degeneracy points of the symbol of the wave operator, in addition to the unique degeneracy point of the problem on the unbounded $\beta$-plane. We will demonstrate that these additional degeneracy points bear topological charges which compensate the $+2$ Chern number of the equatorial degeneracy point, thus making the eigenbands of the symbol have zero total Chern numbers, which justifies the absence of transiting modes and the breaking of Yanai and Kelvin mode branches on the sphere for large $\epsilon$.
@@ 4.1 Hermitian form, analogy with topographic shallow-water waves
In order to apply the index theorem, we first need to express Equations (4) in the form $\mathcal{H}X=\omega X$, where $\mathcal{H}$ is a Hermitian differential wave operator and the complex vector $X$ contains the three dependent variables $\tilde{u}(\theta),\tilde{v}(\theta)$ and $\tilde{h}(\theta)$. To do this, we introduce the rescaled fields $u=\sqrt{\cos(\theta)}\,\tilde{u}$, $v=\sqrt{\cos(\theta)}\,\tilde{v}$ and $\eta=\sqrt{\cos(\theta)}\,\tilde{h}$. Equations (4) can thus be recast in the form of a matrix eigenvalue equation:
$\omega\begin{pmatrix}u\\[6.0pt] v\\[6.0pt] \eta\end{pmatrix}=\begin{pmatrix}0&&{\rm i}f(\theta)&&\frac{c}{R}\frac{m}{\cos\theta}\\[6.0pt] -{\rm i}f(\theta)&&0&&-{\rm i}\frac{c}{R}\frac{{\rm d}}{{\rm d}\theta}-{\rm i}\beta_{\rm g}(\theta)\\[6.0pt] \frac{c}{R}\frac{m}{\cos\theta}&&-{\rm i}\frac{c}{R}\frac{{\rm d}}{{\rm d}\theta}+{\rm i}\beta_{\rm g}(\theta)&&0\end{pmatrix}\begin{pmatrix}u\\[6.0pt] v\\[6.0pt] \eta\end{pmatrix}=\mathcal{H}\begin{pmatrix}u\\[6.0pt] v\\[6.0pt] \eta\end{pmatrix}\ ,$ (11)
@@
where we have also defined $M=c\>m/(R\cos\theta)$, and $K_{\theta}$ the symbol of the Hermitian differential operator $-{\rm i}{\rm d}/{\rm d}\theta$ times $c/R$, which is therefore a real quantity (see [63], Theorem 13.8). Generally speaking, the symbol of a wave operator, which is obtained via the Wigner transform (see appendix 6.3), provides a local representation of the wave equations in phase space (see e.g. [64, 65, 66, 62]). It extends the Fourier transform to systems of wave equations with non constant coefficients. In the present case, the symbol $H$ represents quasi-local dispersion and polarisation relations of shallow-water waves on the sphere. $H$ is a continuous matrix function of the conjugated variables $\theta,K_{\theta}$, and $m$. To be clear, $m$ must be an integer so that the wave functions of $\mathcal{H}$ are regular on the sphere, but we can formally consider the symbol $H$ as a continuous function of $m$ or $M$. In the context of this paper, we introduce it in order to apply the index theorem, following [23], i.e. we investigate the topological properties of the degeneracy points of $H$. This theorem relates the Chern numbers of these degeneracy points to the number and direction of transiting modes of the wave operator $\mathcal{H}$, which can be interpreted as the spectral footprint of a degeneracy point of non-zero Chern number in the quasi-local dispersion relations. The symbol (13) is a generalisation of the one studied by [1] for the unbounded $\beta$-plane: indeed, in the Matsuno limit, $\beta_{\rm g}/f\approx\epsilon/4\ll 1$ in the Yoshida waveguide, thus the metric term $\beta_{\rm g}$ can be dropped, as far as equatorially-trapped waves are concerned.
A recent paper [67] wrongly states that the shallow-water wave operator on the sphere is non-Hermitian and admits complex eigenvalues, which is not true since this system is linearly stable in the absence of a background flow. The mistake originates from using the Fourier transform (which is wrong since the coefficients of $\mathcal{H}$ vary with $\theta$) on the wave operator, instead of the more relevant Wigner transform. It is known that the Wigner transform must include additional terms on curved manifolds [68], which yields a Hermitian symbol in this case. Moreover, the equivalence between the respective Hermiticity of an operator and its symbol is a known property [63], and the symbol (13) is Hermitian as we used an appropriate rescaling of the variables, implying the Hermiticity of the operator $\mathcal{H}$ for the canonical scalar product. Besides, we wish to point out that the topological properties that are presented in the following section do not depend on the choice of $\theta$ (and thus its conjugated symbol $K_{\theta}$) as coordinate to express the differential operator $\mathcal{H}$ and its symbol. Indeed, one would obtain the same symbol $H$ using any alternative coordinate $x=F(\theta)$, as long as the fields are appropriately rescaled so as to preserve the Hermiticity of $\mathcal{H}$ (see appendix 6.3).
We also wish to point out that Equations (11) are formally equivalent to the shallow-water model in planar geometry with varying topography, whose transiting modes were investigated by [27]. The metric term $\beta_{\rm g}$ plays the exact same role as the topographic parameter $\beta_{t}$, although [27] only considered an $f$-plane. In other words, the metric $\beta$-term is mathematically analogous to topography in flat metric. Actually, one could consider together the effect of a curved metric (spherical or oblate for rapidly rotating celestial bodies, for instance) with a topography $h_{0}(\theta)$ varying with latitude (thus the velocity $c=\sqrt{gh_{0}}$ is also a function of latitude). It can be shown that the symbol of this problem is the same as (13), with a total $\beta$-term that is the sum of $\beta_{\rm g}$ and $\beta_{t}$. For the spherical metric with topography, this adds up to $\beta_{\text{total}}=(c/2R)\left(\tan\theta-{\rm d}\ln c/{\rm d}\theta\right)$. We will not be considering a varying topography, as we wish to focus on the curved metric. However, we bring to the analysis of [27] the case of varying $f$ and the additional geometric/topographic term with latitude, which was already addressed by [37] (part 3.5.1), in the context of the equatorial channel.